# InternLM2.5-StepProver: Advancing Automated Theorem Proving via Expert Iteration on Large-Scale LEAN Problems

### v1

@@ 3.1 Results of Expert Iteration
3.2 Benchmark Performance of InternLM2.5-StepProver
We have conducted a comprehensive analysis of InternLM2.5-StepProver on several standard formal benchmarks, in comparison with our previous model InternLM2-StepProver, as well as a number of other frontier language models, to exhibit the strength of our approach.
MiniF2F
We first analyze the performance on the MiniF2F benchmark Zheng et al. (2021). The original benchmark was released in Lean 3 and was later ported to an earlier version of Lean 4. For our analysis, we use the version of MiniF2F in Lean 4, as released by the LeanDojo project Yang et al. (2024), with adaptations to Lean 4.7.0 and corrections of several formalization mistakes.
The best-first-search approach employs an evaluation setting similar to that of InternLM2-StepProver, where the model selects states to expand based on the average log-likelihood of the tactics leading to those states. In contrast, the Critique-Guided (CG) search method involves the policy model selecting states to expand based on an external critic model that grades each state. The search budget for both methods can be universally described as $P\times S\times K$, where $P$ represents the number of passes, $S$ the number of states, and $K$ the maximum number of state expansions, or search iterations. In our context, we set $S=32$ and $K=600$, with the temperature fixed at $T=0.7$. In a BF+CG scenario, we perform half of the passes using the BF method and the other half using the CG method.
The test results are presented in Tab. 3. From the table, we observe that our approach, InternLM2.5-StepProver, achieves an accuracy rate of 69.6% on MiniF2F-Valid and 65.9% on MiniF2F-Test. This represents a significant improvement of 11.4% over InternLM2-StepProver and outperforms all other approaches, including those using formalization languages different from Lean 4. In addition, in a BF scenario, InternLM2.5-StepProver also achieves better inference scalability with a 59.2% pass@64, significantly higher than its previous model.
ProofNet and Putnam
The evaluation setting is mainly the same as MiniF2F, except that due to the restriction of compute resources, we only report the performance of the BF+CG approach here. Tab. 4 exhibits the performance of various models on the ProofNet dataset. We use BF+CG strategy by default, and equally distribute our search budget on two methods. As we are not taking the validation set of ProofNet for expert iteration, we only report the pass rate on the whole dataset. InternLM2.5-StepProver achieved a pass@256 of 27.0% for the overall ProofNet dataset, surpassing the existing state-of-the-art methods, Deepseek-Prover-V1.5-RL (25.3%).
Tab. 5 shows the performance of our approach on the Putnam dataset. We choose $P=2$ in this scenario as our method requires at least two independent passes, i.e. one for Best First and one for Critique Guided. Our approach, without the informal proof skeleton, surpasses all prior methods on the benchmark with 6 problems solved. As the results indicate, the expert iteration, together with the critique model, provides decent performance improvement even on tasks with an unseen distribution.
Method Model size Pass MiniF2F-valid MiniF2F-test
Whole-Proof Generation Methods
TheoremLlama (Wang et al., 2024) - cumulative $36.5\%$ $33.6\%$
DeepSeek-Prover (Xin et al., 2024a) 7B $128$ $46.1\%\pm 0.5\%$
$16\times 4096$ $50.0\%$
DeepSeek-Prover-V1.5-RL 7B 32 - $50.0\%\pm 0.5\%$
64 - $50.7\%\pm 0.4\%$
$128$ $51.6\%\pm 0.5\%$
$3200$ $54.9\%\pm 0.7\%$
$4\times 6400$ $58.4\%\pm 0.6\%$
$16\times 6400$ ${60.2\%}$
Tree Search Methods
PACT (Han et al., 2021) 837M $1\times 16\times 512$ $23.9\%$ $24.6\%$
$8\times 16\times 512$ $29.3\%$ $29.2\%$
ReProver (Yang et al., 2024) 229M - - $26.5\%$
Llemma (Azerbayev et al., 2023b) 7B $1\times 32\times 100$ $26.2\%$ $26.2\%$
$8\times 8\times 512$ $41.2\%$ $34.5\%$
$64\times 8\times 512$ $47.3\%$ $36.6\%$
HTPS (Lample et al., 2022) 600M cumulative $58.6\%$
$64\times 5000$ $41.0\%$
Lean-STaR  (Lin et al., 2024) 7B $64\times 1\times 50$ $46.3\%$
InternLM2-Math-Plus 7B $1\times 32\times 100$ $43.4\%$
DeepSeek-Prover-V1.5-RL 7B $1\times 3200$ $55.0\%\pm 0.7\%$
$4\times 6400$ $59.6\%\pm 0.6\%$
$16\times 6400$ ${62.7\%}$
$32\times 6400$ ${63.5\%}$
$64\times 32\times 100$ ${63.9\%}$ ${54.5\%}$
InternLM2.5-StepProver-BF 7B $1\times 32\times 600$ $55.4\%\pm 1.5\%$ $47.3\%\pm 1.1\%$
$4\times 32\times 600$ ${61.3\%\pm 0.7\%}$ ${52.6\%}\pm 1.0\%$
$16\times 32\times 600$ ${63.7\%\pm 0.4\%}$ ${57.3\%}\pm 0.7\%$
$64\times 32\times 600$ ${64.6\%\pm 0.3\%}$ ${59.2\%}\pm 0.1\%$
$256\times 32\times 600$ ${65.1\%}$ ${59.4\%}$
InternLM2.5-StepProver-BF+CG 7B $2\times 32\times 600$ $56.0\%\pm 1.5\%$ $50.7\%\pm 1.3\%$
$4\times 32\times 600$ ${61.4\%\pm 0.6\%}$ ${58.5\%}\pm 0.9\%$
$16\times 32\times 600$ ${65.8\%\pm 0.5\%}$ ${62.5\%}\pm 0.5\%$
$64\times 32\times 600$ ${68.0\%\pm 0.3\%}$ ${63.8\%}\pm 0.3\%$
$256\times 32\times 600$ 69.6% 65.9%
Table 4: Comparing with state-of-the-art on the ProofNet (Azerbayev et al., 2023a) dataset.
Method Pass ProofNet
valid test all
ReProver (Yang et al., 2024) - - - $13.8\%$
Deepseek-Prover-V1.5-RL $1\times 3200$ $22.0\%\pm 0.3\%$ $21.5\%\pm 0.8\%$ $21.8\%\pm 0.4\%$
$4\times 6400$ $25.4\%$ $25.3\%$ $25.3\%$
InternLM2-StepProver $1\times 32\times 100$ $18.1\%$
InternLM2.5-StepProver-BF+CG $4\times 32\times 600$ $18.8\%\pm 0.7\%$
$64\times 32\times 600$ $23.6\%\pm 0.4\%$
$256\times 32\times 600$ $27.0\%$
Table 5: Comparing with state-of-the-art on the Putnam (Tsoukalas et al., 2024) dataset.
Method Model size Pass Result
GPT-4 (Achiam et al., 2023) - $01$/640
COPRA (GPT-4) (Thakur et al., 2023) - $01$/640
DSP(Isabelle) (Jiang et al., 2022) 540B $10$ 4/640
ReProver (Yang et al., 2024) 229M $1$ 0/640
InternLM2-StepProver 7B 1 5/640
InternLM2.5-StepProver-BF+CG 7B 2 6/640
@@ 4 Related Work
Automated Theorem Proving does not have a unified approach. The mainstream paradigm is training a language model on tuples of (proof state, next tactic), followed by a tree search to find proofs (Polu & Sutskever, 2020; Polu et al., 2022; Lample et al., 2022; Yang et al., 2024; Lin et al., 2024; Wu et al., 2024). Another line of the work is training to auto-regressive generate a whole proof based on a theorem statement from the model itself (Xin et al., 2024a; b) or translating from human informal proof (Jiang et al., 2022; Wu et al., 2022; Wang et al., 2024). No matter learning with which paradigms, most methods rely on expert iteration (Anthony et al., 2017) for model self-improving. In this work, we train our model with a state-tactic paradigm via expert iteration on a large-scale LEAN dataset.
@@ 5 Conclusion
In this paper, we introduce InternLM2.5-StepProver which improves its automated theorem-proving ability via large-scale expert iteration and achieves state-of-the-art on multiple benchmarks. We use over 20,000 CPU days to search for proofs on Lean-Workbook-Plus and prove or disprove 17% of them. We analyze the difficult distribution of Lean-Workbook-Plus based on the proof lengths and CPU usage. We observe a log-linear trend between proof length/CPU usage with proof problem amount. To prove more problems in Lean-Workbook-Plus, we need to put more effort into search budgets and improve the search algorithm efficacy. Our findings provide concrete guidance for future developments in automated mathematical reasoning.
@@
This work is mainly focused on context-level math problems and pays less attention to other automated theorem-proving scenarios. We do not have a stable metric to measure critic models which makes iteration of critic models difficult.
@@ Appendix A Training details
Policy Model
InternLM2.5-StepProver’s policy model is built upon InternLM-math-plus-7B Ying et al. (2024b). We used the same training setting when we performed the expert iteration process: We used a global batch size of 512 and a learning rate of $2\times 10^{-5}$. We fine-tuned for 2 epochs to obtain the SFT model. For the learning rate, we used a warm-up in the first 3% steps, followed by a cosine schedule decaying to zero. The entire expert iteration process generated 2.19 billion tokens of data, with the final iteration taking approximately 14 hours on 32 A800 GPUs.
@@ Appendix B Case studies
Case: MiniF2F: mathd_algebra_31 Natural Language problem: If $\sqrt{x+\!\sqrt{x+\!\sqrt{x+\!\sqrt{x+\cdots}}}}=9$, find $x$. Show that it is 72. theorem mathd_algebra_31 (x : NNReal) (u : NNReal) (h : n, u (n + 1) = NNReal.sqrt (x + u n)) (h : Filter.Tendsto u Filter.atTop ( 9)) : 9 = NNReal.sqrt (x + 9) := by have h := h₁.const_add x have h : Filter.Tendsto (fun k => NNReal.sqrt (x + u k)) Filter.atTop ( (NNReal.sqrt (x + 9))) := NNReal.continuous_sqrt.continuousAt.tendsto.comp h have h : (fun k : => NNReal.sqrt (x + u k)) = fun k : => u (k + 1) := by ext k rw [h₀] have h : Filter.Tendsto (fun k : => u (k + 1)) Filter.atTop ( (NNReal.sqrt (x + 9))) := h h have h : Filter.Tendsto (fun k => u (k + 1)) Filter.atTop ( 9) := h₁.comp (Filter.tendsto_add_atTop_nat 1) exact tendsto_nhds_unique h h
It is an interesting case to show how InternLM2.5-StepProver solves problem whose formalized version is significantly harder than the informal one. The informal solution of this problem is not rigorous, which jumps from the equation $\sqrt{x+\!\sqrt{x+\!\sqrt{x+\!\sqrt{x+\cdots}}}}=9$ to $\sqrt{x+9}=9$, involving a substitution that is intuitional but risky. The formalized version of this problem uses series and limitations to redefine the problem. That is the case when formal reasoning is detached from informal reasoning. In such cases, it is hard to say that the involvement of informal CoT has any benefit to the problem. InternLM2.5-StepProver provides a solid proof of the problem without the augmentation of informal information, successfully solving the problem.
Case: ProofNet: Munkers_31_2 Natural Language problem: Show that if $X$ is normal, every pair of disjoint closed sets have neighborhoods whose closures are disjoint. theorem exercise_Munkers_31_2 {X : Type*} [TopologicalSpace X] [NormalSpace X] {A B : Set X} (hA : IsClosed A) (hB : IsClosed B) (hAB : Disjoint A B) : (U V : Set X), IsOpen U IsOpen V A U B V closure U closure V = := by obtain U₀, V₀, hU₀, hV₀, hA₀, hB₀, hAB₀⟩ := normal_separation hA hB hAB obtain U, hU₁, hU₂, hU₃⟩ := normal_exists_closure_subset hA hU hA obtain V, hV₁, hV₂, hV₃⟩ := normal_exists_closure_subset hB hV hB refine U, V, hU₁, hV₁, hU₂, hV₂,?_ exact (hAB₀.mono hU hV₃).eq_bot
InternLM2.5-StepProver has an improved capability of solving undergraduate problems, even if it is not fine-tuned on such data distributions. InternLM2.5-StepProver utilizes premises from the Mathlib to construct valid closure subsets, effectively showing its math reasoning ability.
Appendix C Model Chat Templates
Here are example inputs of our policy and critic models.
An example of the policy model Prompt NAME: square_sub_one_divisible_eight PROOF_BEFORE: rw [h, pow_two] STATE_BEFORE: m n : h : n = 2 * m + 1 8 | (2 * m + 1) * (2 * m + 1) - 1 TACTIC: Response rw [← Nat.mod_add_div (2 * m + 1) 8]
[…]

### v2

@@ 4 Related Work
Automatic Theorem Proving.
Automated theorem proving does not have a unified approach. A mainstream paradigm is training a language model on tuples of (proof state, next tactic), followed by a tree search to find proofs (Polu & Sutskever, 2020; Polu et al., 2022; Lample et al., 2022; Yang et al., 2024; Lin et al., 2024; Wu et al., 2024). Another line of work is training to auto-regressively generate a whole proof based on a theorem statement from the model itself (Xin et al., 2024a, b) or translating from human informal proofs (Jiang et al., 2022; Wu et al., 2022; Wang et al., 2024b). Regardless of the learning paradigm used, most methods rely on expert iteration (Anthony et al., 2017) to improve the model and require the prover model itself to determine the search direction. In this work, we train our prover model to search under the guidance of an external critic model. The models are trained with a state-tactic paradigm via expert iteration on a large-scale LEAN dataset.
Process-supervised Reasoning in Informal Mathematics.
Recently, numerous studies have shown that allowing an external language model to evaluate the quality of reasoning steps can boost performance on tasks such as informal math (Lightman et al., 2023; Zheng et al., 2025; Uesato et al., 2022). Although this approach has proven effective, it often requires a significant amount of human-annotated data or extensive computation (Luo et al., 2024; Wang et al., 2024a). Furthermore, such methods are restricted to informal language reasoning tasks. We propose a new scheme for training a tactic-level critic model that leverages the correctness signal from the formal system.
@@ 5 Conclusion
In this paper, we presented critic-guided search, a novel approach that significantly enhances the theorem proving capabilities of prover language models in formal mathematics by using an external critic model to rerank intermediate Lean states. Our method begins with training initial critic models from the trajectories of prior prover models to predict preference information between Lean states. We further improved the prover and the critic in tandem using expert iteration, fine-tuning the models on correct proof trajectories that the prover model samples and verifies using the Lean solver. Our contributions include the introduction of the theorem proving dataset Lean-Workbook-Plus, demonstrating that critic-guided search can further improve search depth, and achieving new results on the miniF2F-test benchmark, increasing the pass rate from 54.5% to 65.9%.
These advancements not only improve the performance of automated theorem proving, but also offer concrete guidance for future developments in automated mathematical reasoning.
@@
This work is mainly focused on contest-level math problems and pays less attention to other automated theorem-proving scenarios. Besides, we currently do not have a stable metric to measure critic models, which makes iteration of critic models difficult.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Training details
Prover Model
Our prover model is built upon InternLM-math-plus-7B (Ying et al., 2024b). We used the same training setting when we performed the expert iteration process: We used a global batch size of 512 and a learning rate of $2\times 10^{-5}$. We fine-tuned for 2 epochs to obtain the SFT model. For the learning rate, we used a warm-up in the first 3% steps, followed by a cosine schedule decaying to zero. The entire expert iteration process generated 2.19 billion tokens of data, with the final iteration taking approximately 14 hours on 32 A800 GPUs.
@@ Appendix B Case studies
Case: miniF2F: mathd_algebra_31 Natural Language problem: If $\sqrt{x+\!\sqrt{x+\!\sqrt{x+\!\sqrt{x+\cdots}}}}=9$, find $x$. Show that it is 72. theorem mathd_algebra_31 (x : NNReal) (u : NNReal) (h : n, u (n + 1) = NNReal.sqrt (x + u n)) (h : Filter.Tendsto u Filter.atTop (mathscrN 9)) : 9 = NNReal.sqrt (x + 9) := by have h := h₁.const_add x have h : Filter.Tendsto (fun k => NNReal.sqrt (x + u k)) Filter.atTop (mathscrN (NNReal.sqrt (x + 9))) := NNReal.continuous_sqrt.continuousAt.tendsto.comp h have h : (fun k : => NNReal.sqrt (x + u k)) = fun k : => u (k + 1) := by ext k rw [h₀] have h : Filter.Tendsto (fun k : => u (k + 1)) Filter.atTop (mathscrN (NNReal.sqrt (x + 9))) := h h have h : Filter.Tendsto (fun k => u (k + 1)) Filter.atTop (mathscrN 9) := h₁.comp (Filter.tendsto_add_atTop_nat 1) exact tendsto_nhds_unique h h
This case demonstrates how InternLM2.5-StepProver solves a problem whose formalized version is significantly harder than the informal one. The informal solution of this problem is not rigorous, which jumps from the equation $\sqrt{x+\!\sqrt{x+\!\sqrt{x+\!\sqrt{x+\cdots}}}}=9$ to $\sqrt{x+9}=9$, involving a substitution that is intuitive but risky. The formalized version of this problem uses series and limitations to redefine the problem. This is a case where formal reasoning detaches from informal reasoning. In such cases, it is hard to say that the involvement of informal CoT has any benefit to the problem. InternLM2.5-StepProver provides a solid proof of the problem without the augmentation of informal information, successfully solving the problem.
Case: ProofNet: Munkers_31_2 Natural Language problem: Show that if $X$ is normal, every pair of disjoint closed sets has neighborhoods whose closures are disjoint. theorem exercise_Munkers_31_2 {X : Type*} [TopologicalSpace X] [NormalSpace X] {A B : Set X} (hA : IsClosed A) (hB : IsClosed B) (hAB : Disjoint A B) : (U V : Set X), IsOpen U IsOpen V A U B V closure U closure V = := by obtain U₀, V₀, hU₀, hV₀, hA₀, hB₀, hAB₀⟩ := normal_separation hA hB hAB obtain U, hU₁, hU₂, hU₃⟩ := normal_exists_closure_subset hA hU hA obtain V, hV₁, hV₂, hV₃⟩ := normal_exists_closure_subset hB hV hB refine U, V, hU₁, hV₁, hU₂, hV₂,?_ exact (hAB₀.mono hU hV₃).eq_bot
InternLM2.5-StepProver has an improved capability of solving undergraduate problems, even though it was not fine-tuned on such data distributions. InternLM2.5-StepProver utilizes premises from the Mathlib to construct valid closure subsets, effectively demonstrating its mathematical reasoning ability.